OpenStudy (anonymous):

Write the equation of a circle with center at (0, 0) and a diameter of 20.

OpenStudy (anonymous):

x^2+y^2=r^2

OpenStudy (ghazi):

(x-0)^2+(y-0)^2=(10)^2

OpenStudy (ghazi):

equation of circle is\[(x-h)^2+(y-k)^2=(r)^2\] h,k center and r-- radius

OpenStudy (ghazi):

dia=20 radius= dia/2=20/2=10 :)